### v1

@@ 6.2 X-ray emitting phases in numerical simulations
To compare the measured physical quantities with those in numerical simulations, we use the public data (Nelson et al. 2019) of IllustrisTNG simulations (Nelson et al. 2018; Pillepich et al. 2018; Naiman et al. 2018; Springel et al. 2018; Marinacci et al. 2018) and select the $z=0.48$ snapshot of the TNG300-1 box. We aim to create phase diagrams of the filament gas. We use the DisPerSE algorithm to construct the filaments in that snapshot using $129<\log(M_{*}/M_{\sun})<$ 12 galaxies (Nelson et al. 2019). The detailed procedure follows Galárraga-Espinosa et al. (2021) and Malavasi et al. (2022). The calculations of the distances of each gas cell to the nearest filament $d_{\mathrm{fil}}$, to the nearest maxima or bifurcation critical point $d_{\mathrm{CP}}$, and the nearest critical point but along the direction of filament $d_{\mathrm{skel}}$ follow the definition in Malavasi et al. (2022).
@@
To reduce the computational time, we randomly select 1/500 of the total gas cells, which is enough to represent the phase diagram (Galárraga-Espinosa et al. 2021). We select gas cells with $d_{\mathrm{fil}}<1$ Mpc and $d_{\mathrm{CP}}>2$ Mpc, and remove cells in halos or within a 25 Mpc range to simulation box boundaries. The temperature of a gas cell is calculated by using the quantities InternalEnergy and ElectronAbundance based on the standard recipe [^7] https://www.tng-project.org/data/docs/faq/#gen6. The baryon density contrast of a cell is converted from the density by comparing it with the mean baryon density at the redshift of 0.48. We plot the mass-weighted phase diagrams of filament gas in the left panel of Fig. 10, where the majority of the gas is in $10^{4.5}-10^{6.5}$ K temperature and $1-100$ baryon density contrast phases. We note that the presence of hot phases with $T>10^{7}$ K is not as prominent as those in the $z=0$ snapshot (Galárraga-Espinosa et al. 2021). It could be due to the redshift evolution of filament gas, which is being continuously heated during structure formation (Cui et al. 2019; Martizzi et al. 2019).
@@
To investigate the traceable populations of the filament gas, we create two phase diagrams weighted by electron pressure and $0.5-2$ keV photon luminosity folded by eROSITA response, respectively (middle and right panels in Fig. 10). The former one represents the traceable population of SZ observations (e.g. de Graaff et al. 2019; Tanimura et al. 2019, 2020a), and the latter one represents the traceable population of soft band X-ray emission studies. The two diagrams show that the traceable phases of SZ signal and soft band X-ray emission are different. SZ observations trace $10^{5.5}-10^{6.8}$ K temperature and a $\Delta_{\mathrm{b}}=10^{0.3}-10^{2}$ density phases WHIM in filaments.On the other hand, X-ray emission in the 0.5–2 keV band traces $10^{6.2}-10^{6.9}$ K temperature and $\Delta_{\mathrm{b}}=10^{1.2}-10^{2.2}$ density phases WHIM. It also shows that the X-ray emitting population only contributes a fraction of the total SZ signal, which agrees with the assumption we used in Sect. 6.1. Our averaged density contrast measurement agrees with the TNG numerical simulations. However, the best-fit temperature using a single component model is close to the upper boundary of the contour of X-ray emitting WHIM (see the yellow data point with error bar in the left panel of Fig.10). This discrepancy has been addressed in Sect. 5.2, i.e., the best-fit temperature value using a 1-T model will be biased to the high end of the temperature distribution when fitting a spectrum with multi-temperature nature.
@@ 6.3 Impact of photoionization
It is known that the WHIM is also partially ionized by the cosmic cosmic ultraviolet background (UVB) (e.g., Nicastro et al. 1999; Krongold et al. 2003), which alters the ionization balance and exhibit different spectral features from the CIE plasma (Khabibullin & Churazov 2019; Churazov et al. 2023). In this section, we combine the WHIM phase distributions in numerical simulations (left panel of Fig. 10) and the photoionization emission model to characterize the impact of including photoionization model on our conclusions.
@@
where $F_{\mathrm{1-1000Ry}}$ is the local flux in the 1–1000 Ry energy band (1 Ry = 13.6 eV) following the convention of Krolik et al. (1981). We adopt the Faucher-Giguère (2020) UVB spectral energy distribution for the calculation. We obtain a relation of $\log\xi[\mathrm{erg\ s^{-1}\ cm^{2}}]=2.26-\log\Delta_{\mathrm{b}}$ at $z=0.48$. We use the pion model (Miller et al. 2015; Mehdipour et al. 2016) in SPEX-3.08.00 (Kaastra et al. 1996, 2024) for calculating the photoioized emission spectra and energy loss rates due to radiation at different $\xi$ and temperature. We note that by setting the parameter tmod=1, the pion model in SPEX allows calculations at fixed temperatures and ionization parameters without solving the energy balance, which meets the physical condition of WHIM, whose temperature is higher than the photoionization equilibrium temperature and is maintained by external heating mechanisms such as shock heating.
Figure 11: Top: Photon luminosity per unit of emission measure in the 0.5–2 keV band of the photoionized WHIM at different density (x-axis) and temperature (y-axis) by assuming $Z=0.2Z_{\sun}$. Towards the high density (low ionization parameter) end, the radiative cooling reaches the CIE limit and is simply a function of temperature. When $\Delta_{\mathrm{b}}\lesssim 1$, the radiative cooling keeps high at low temperature. Bottom: Same as the right panel of Fig. 10 but calculated using the photoionization model. By taking the cosmic UV background photoionization into account, the size of the 50% enclosed phases contour is marginally larger than that of the CIE condition.
The top panel of the Fig. 11 shows the 0.5–2 keV photon luminosity per unit of emission measure of the UVB photonionized WHIM at different temperature and density contrast. When $\Delta_{\mathrm{b}}>2$ ($\log\xi[\mathrm{erg\ s^{-1}\ cm^{2}}]\lesssim 0)$, the radiative cooling reaches the CIE limit and is simply a function of temperature. When $\Delta_{\mathrm{b}}\lesssim 1$ ($\log\xi[\mathrm{erg\ s^{-1}\ cm^{2}}]\gtrsim 1)$, the radiative cooling deviates from the CIE condition and is a function of both temperature and density. The radiative cooling in the 0.5–2 keV band keeps high for the WHIM with temperature below $10^{6}$ K. We use the results of the photon luminosity of the photoionized WHIM to update the phase diagram weighted by the 0.5–2 keV emission and plot it in the bottom panel of Fig. 11. The 50% enclosed phases contour is at the same position of that of the CIE condition with a slightly larger size. It means that our conclusion drawn from Sect. 6 does not change by taking photonionization into account. The traceable WHIM by broadband X-ray emission are in the highest temperature and density phases among the whole filament gas population.
@@ 7 Summary
eRASS provides us with the deepest All-Sky X-ray survey in the soft band, allowing a more precise investigation of X-ray emission in cosmic filaments than previous pioneering works (Tanimura et al. 2020b, 2022). The aim of this work was to use the eRASS:4 X-ray data and SDSS optical filament catalog to stack X-ray emission from filaments and study the properties of the X-ray emitting WHIM. Comparing to the pioneering works, we adopt a novel method to estimate the fraction of non-WHIM emission in the stacked signal; improve the spectral stacking method to minimize the systematics from the instrumental and sky background features at the observer frame; combine the SZ stacking results to quantify the volume fraction and density of the X-ray emitting WHIM simultaneously. We summarize our main results as follows:
@@
    We stacked 0.3–1.2 keV surface brightness profiles of 7817 SDSS optical filaments with a physical lengths of $20-100$ Mpc and redshifts of 0.2–0.6 in a 2275 $\deg^{2}$ footprint and detect an X-ray emission excess over the background at a $9\sigma$ significance.
@@
    We validated our profile stacking method by applying it to simulated count maps that include emission from filaments with an injected model and randomized X-ray background based on the observed angular power spectrum. After 200 realizations, the recovered averaged filament profiles agree well with the injected model with only 13% difference, which could be from the possible boosting from projectively overlapped filaments in same orientations.
@@
    We carefully estimate the fraction of stacked signal from unmasked sources including X-ray halos, AGNs and XRBs that is associated with galaxies. We first paint the observed galaxies in the analysis footprint with X-ray count rate from $L_{\mathrm{X}}-M_{*}$ scaling relations to create an expected X-ray emission map from galaxies, where the incompleteness and bias of observed galaxy SMF is corrected. Then we apply the same surface brightness profile stacking method to obtain the fractions of stacked signal from each types of unmasked sources, which is 20%, 12%, and 5% for X-ray halos, AGNs, and XRBs, respectively. The photometric redshift uncertainties of the galaxies we used could underestimate another 3% of the contamination fraction. Based on these fractions, we conclude that the remaining $\sim 60\%$ of the signal is from WHIM in the filaments and the WHIM detection significance is $5.4\sigma$.
@@
    We stacked broadband spectra and response files of the filaments in the rest-frame. We included the components of unmasked sources in spectral fitting. The best-fit temperature by using a 1-T model is $0.58\pm 0.10$ keV, which could be biased to the high end of the WHIM temperature distribution given its multi-temperature nature suggested by numerical simulations. The stacked spectra can also be fitted by a numerical simulation motivated model with a log-normal temperature distribution, which is able to reduce the residual in the 0.5–0.6 keV energy range, suggesting the presence of the OVII abundant temperature phases.
@@
    We modeled the baryon density contrast $\Delta_{\mathrm{b}}$ of the X-ray emitting WHIM by fitting the stacked surface brightness profile, where we additionally introduced a parameter of volume filling factor $f_{\mathrm{vol}}$, which is necessary because only a small fraction of WHIM emits X-rays that can be observed by eROSITA. We use the SZ stacking results in literature to break the degeneracy between the $\Delta_{\mathrm{b}}$ and volume fraction $f_{\mathrm{vol}}$, and obtained an averaged $\log\Delta_{\mathrm{b}}=1.88\pm 0.18$, suggesting that the observed emission is from the densest population of the WHIM.
@@
    We analyzed the $z=0.48$ snapshot of the TNG300-1 simulation to investigate the X-ray emitting phases in simulations. We found that most of the photons in the 0.5–2 keV energy range is emitted by phases of $6.96.2<\log(T/\mathrm{K})<$ 6.9 and $2.21.2<\log\Delta_{\mathrm{b}}<$ 2.2, which generally agree with our estimated values.
It is the first time we detect $>5\sigma$ X-ray emission from cosmic filaments and apply a thorough investigation of the properties of the X-ray emitting WHIM. In the coming decade, our understanding of X-ray emissions from filaments, and from the WHIM within them, will be significantly improved from multiple aspects. The 4MOST Cosmology Redshift Survey (Richard et al. 2019) will bring us a deep and wide field galaxy spectroscopic redshift data in the southern hemisphere, allowing us to take full use of the eRASS data of the German consortium. Improved filament finders such as Monte Carlo Physarum Machine (Elek et al. 2022) will be worked on the new survey data to provide a more accurately constructed filament catalog. Meanwhile, better understandings of X-ray properties of galaxy groups, AGNs, and XRBs will help us to better estimate the contamination fractions of unmasked sources. Moreover, high spectral resolution X-ray missions such as HUBS (Cui et al. 2020) and LEM (Kraft et al. 2022) will be able to explore a wider parameter space of the WHIM properties.
@@ Acknowledgements.
We acknowledge Jelle Kaastra for the instruction on using the SPEX Pion model. X.Z., E.B., V.G., A.L., C.G., and S.Z. acknowledge financial support from the European Research Council (ERC) Consolidator Grant under the European Union’s Horizon 2020 research and innovation program (grant agreement CoG DarkQuest No 101002585). M.C.H.Y. and M.F. acknowledge support from the Deutsche Forschungsgemeinschaft through the grant FR 1691/2-1. Y.Z. and G.P. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 865637). G.P. acknowledges support from Bando per il Finanziamento della Ricerca Fondamentale 2022 dell’Istituto Nazionale di Astrofisica (INAF): GO Large program and from the Framework per l’Attrazione e il Rafforzamento delle Eccellenze (FARE) per la ricerca in Italia (R20L5S39T9). M.B. acknowledges funding by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. A.V. acknowledges funding by the Deutsche Forschungsgemeinschaft – 450861021. N. M. acknowledges funding by the European Union through a Marie Skłodowska-Curie Action Postdoctoral Fellowship (Grant Agreement: 101061448, project: MEMORY). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or of the Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument were led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA.

### v2

@@ 6.2 X-ray emitting phases in numerical simulations
To compare the measured physical quantities with those in numerical simulations, we used the public data (Nelson et al. 2019) from IllustrisTNG simulations (Nelson et al. 2018; Pillepich et al. 2018; Naiman et al. 2018; Springel et al. 2018; Marinacci et al. 2018). We selected the $z=0.48$ snapshot of the TNG300-1 box and aimed to create phase diagrams of the filament gas. We used the DisPerSE algorithm to construct the filaments in that snapshot using $129<\log(M_{*}/M_{\sun})<$ 12 galaxies (Nelson et al. 2019). The detailed procedure follows Galárraga-Espinosa et al. (2021) and Malavasi et al. (2022). The calculations of the distances of each gas cell to the nearest filament, $d_{\mathrm{fil}}$, to the nearest maxima or bifurcation critical point, $d_{\mathrm{CP}}$, and the nearest critical point but along the direction of filament, $d_{\mathrm{skel}}$, follow the definition in Malavasi et al. (2022).
@@
To reduce the computational time, we randomly selected 1/500 of the total gas cells, which is enough to represent the phase diagram (Galárraga-Espinosa et al. 2021). We selected gas cells with $d_{\mathrm{fil}}<1$ Mpc and $d_{\mathrm{CP}}>2$ Mpc. Then, we removed cells in halos or within a 25 Mpc range to simulation box boundaries. The temperature of a gas cell is calculated by using the quantities InternalEnergy and ElectronAbundance based on the standard recipe [^7] https://www.tng-project.org/data/docs/faq/#gen6. The baryon density contrast of a cell is converted from the density by comparing it with the mean baryon density at the redshift of 0.48.
We plot the mass-weighted phase diagrams of filament gas in the left panel of Fig. 10, where the majority of the gas is in $10^{4.5}-10^{6.5}$ K temperature and $1-100$ baryon density contrast phases. We note that the presence of hot phases with $T>10^{7}$ K is not as prominent as those in the $z=0$ snapshot (Galárraga-Espinosa et al. 2021). It could be due to the redshift evolution of filament gas, which is being continuously heated during structure formation (Cui et al. 2019; Martizzi et al. 2019).
@@
To investigate the traceable populations of the filament gas, we created two phase diagrams weighted by electron pressure and $0.5-2$ keV photon luminosity folded by eROSITA response, respectively (middle and right panels in Fig. 10). The former one represents the traceable population of SZ observations (e.g. de Graaff et al. 2019; Tanimura et al. 2019, 2020a), while the latter one represents the traceable population of soft band X-ray emission studies. The two diagrams show that the traceable phases of SZ signal and soft band X-ray emission differ. SZ observations trace $10^{5.5}-10^{6.8}$ K temperature and a $\Delta_{\mathrm{b}}=10^{0.3}-10^{2}$ density phases WHIM in filaments.On the other hand, X-ray emission in the 0.5–2 keV band traces $10^{6.2}-10^{6.9}$ K temperature and $\Delta_{\mathrm{b}}=10^{1.2}-10^{2.2}$ density phases WHIM. It also shows that the X-ray emitting population only contributes a fraction of the total SZ signal, which agrees with the assumption we used in Sect. 6.1. Our averaged density contrast measurement agrees with the TNG numerical simulations. However, the best-fit temperature using a single component model is close to the upper boundary of the contour of X-ray emitting WHIM (see the yellow data point with error bar in the left panel of Fig.10). This discrepancy has been addressed in Sect. 5.2, i.e., the best-fit temperature value using a 1-T model will be biased to the high end of the temperature distribution when fitting a spectrum with multi-temperature nature.
@@ 6.3 Impact of photoionization
It is known that the WHIM is also partially ionized by the cosmic cosmic ultraviolet background (UVB), which alters the ionization balance and exhibits different spectral features from the CIE plasma (e.g., Nicastro et al. 1999; Krongold et al. 2003; Khabibullin & Churazov 2019; Churazov et al. 2023). In this section, we describe how we combined the WHIM phase distributions in numerical simulations (left panel of Fig. 10) and the photoionization emission model to characterize the impact of including the photoionization model on our results.
@@
where $F_{\mathrm{1-1000Ry}}$ is the local flux in the 1–1000 Ry energy band (1 Ry = 13.6 eV) following the convention of Krolik et al. (1981). We adopted the Faucher-Giguère (2020) UVB spectral energy distribution for the calculation. We obtained a relation of $\log\xi[\mathrm{erg\ s^{-1}\ cm^{2}}]=2.26-\log\Delta_{\mathrm{b}}$ at $z=0.48$. We used the pion model (Miller et al. 2015; Mehdipour et al. 2016) in SPEX-3.08.00 (Kaastra et al. 1996, 2024) for calculating the photoioized emission spectra and energy loss rates due to radiation at different $\xi$ and temperature. We note that by setting the parameter tmod=1, the pion model in SPEX allows for calculations to be carried out at fixed temperatures and ionization parameters without solving the energy balance, which meets the physical condition of WHIM, whose temperature is higher than the photoionization equilibrium temperature and is maintained by external heating mechanisms such as shock heating.
Figure 11: Impact of the cosmic UV background photoionization on the X-ray detectable phases. Top: Photon luminosity per unit of emission measure in the 0.5–2 keV band of the photoionized WHIM at different density (x-axis) and temperature (y-axis) by assuming $Z=0.2Z_{\sun}$. Towards the high density (low ionization parameter) end, the radiative cooling reaches the CIE limit and is simply a function of temperature. When $\Delta_{\mathrm{b}}\lesssim 1$, the radiative cooling keeps high at low temperature. Bottom: Same as the right panel of Fig. 10 but calculated using the photoionization model. By taking the cosmic UV background photoionization into account, the size of the 50% enclosed phases contour is marginally larger than that of the CIE condition.
The top panel of the Fig. 11 shows the 0.5–2 keV photon luminosity per unit of emission measure of the UVB photonionized WHIM at different temperature and density contrast. When $\Delta_{\mathrm{b}}>2$ ($\log\xi[\mathrm{erg\ s^{-1}\ cm^{2}}]\lesssim 0)$, the radiative cooling reaches the CIE limit and simply becomes a function of temperature. When $\Delta_{\mathrm{b}}\lesssim 1$ ($\log\xi[\mathrm{erg\ s^{-1}\ cm^{2}}]\gtrsim 1)$, the radiative cooling deviates from the CIE condition and is a function of both temperature and density. The radiative cooling in the 0.5–2 keV band remains high for the WHIM, with a temperature below $10^{6}$ K. We use the results of the photon luminosity of the photoionized WHIM to update the phase diagram weighted by the 0.5–2 keV emission and plot it in the bottom panel of Fig. 11. The 50% enclosed phases contour is at the same position as that of the CIE condition with a slightly larger size. It means that our conclusion drawn from Sect. 6 does not change by taking photonionization into account. The traceable WHIM by broadband X-ray emission is seen to be in the highest temperature and density phases among the whole filament gas population.
@@ 7 Summary
Overall, eRASS has provided the deepest All-Sky X-ray survey in the soft band, allowing for a more precise investigation of X-ray emission in cosmic filaments than in previous pioneering works (Tanimura et al. 2020b, 2022). This work aims to use the eRASS:4 X-ray data and SDSS optical filament catalog to stack X-ray emission from filaments and study the properties of the X-ray emitting WHIM. We adopted a novel method to: 1) estimate the fraction of non-WHIM emission in the stacked signal; 2) improve the spectral stacking method to minimize the systematics from the instrumental and sky background features at the observer frame; and 3) combine the SZ stacking results to simultaneously quantify the volume fraction and density of the X-ray emitting WHIM. We summarize our main results as follows:
@@
    We stacked 0.3–1.2 keV surface brightness profiles of 7817 SDSS optical filaments with physical lengths of $20-100$ Mpc and redshifts of 0.2–0.6 in a 2275 $\deg^{2}$ footprint. We detected an X-ray emission excess over the background at a $9\sigma$ significance.
@@
    We validated our profile stacking method by applying it to simulated count maps, including emission from filaments with an injected model and randomized X-ray background based on the observed angular power spectrum. After 200 realizations, the recovered averaged filament profiles agree well with the injected model with only a 13% difference, which could be the result of a possible boosting from projectively overlapped filaments in the same orientations.
@@
    We carefully estimate the fraction of stacked signal from unmasked sources, including X-ray halos, AGNs, and XRBs that are associated with galaxies. We first paint the observed galaxies in the analysis footprint with X-ray count rate from $L_{\mathrm{X}}-M_{*}$ scaling relations to create an expected X-ray emission map from galaxies, where the incompleteness and bias of observed galaxy SMF is corrected. Then we apply the same surface brightness profile stacking method to obtain the fractions of stacked signal from each type of unmasked sources, which is 20%, 12%, and 5% for X-ray halos, AGNs, and XRBs, respectively. The photometric redshift uncertainties of the galaxies we used could underestimate another 3% of the contamination fraction. Based on these fractions, we conclude that the remaining $\sim 60\%$ of the signal is from WHIM in the filaments and the WHIM detection significance is $5.4\sigma$.
@@
    We stacked broadband spectra and response files of the filaments in the rest frame. We included the components of unmasked sources in spectral fitting. The best-fit temperature using a 1-T model is $0.58\pm 0.10$ keV, which could be biased to the high end of the WHIM temperature distribution given its multi-temperature nature suggested by numerical simulations. The stacked spectra can also be fitted by a numerical simulation-motivated model with a log-normal temperature distribution, which will reduce the residual in the 0.5–0.6 keV energy range, suggesting the presence of the OVII abundant temperature phases.
@@
    We modeled the baryon density contrast $\Delta_{\mathrm{b}}$ of the X-ray emitting WHIM by fitting the stacked surface brightness profile, where we additionally introduced a parameter of volume filling factor $f_{\mathrm{vol}}$, which is necessary because only a small fraction of WHIM emits X-rays that can be observed by eROSITA. We used the SZ stacking results in the literature to break the degeneracy between the $\Delta_{\mathrm{b}}$ and volume fraction $f_{\mathrm{vol}}$ and obtained an averaged $\log\Delta_{\mathrm{b}}=1.88\pm 0.18$, suggesting that the observed emission is from the densest population of the WHIM.
@@
    We analyzed the $z=0.48$ snapshot of the TNG300-1 simulation to investigate the X-ray emitting phases in simulations. We found that most of the photons in the 0.5–2 keV energy range are emitted by phases of $6.96.2<\log(T/\mathrm{K})<$ 6.9 and $2.21.2<\log\Delta_{\mathrm{b}}<$ 2.2, which is in general agreement with our estimated values.
This is the first time we detected $>5\sigma$ X-ray emission from cosmic filaments. We applied a thorough investigation of the properties of the X-ray emitting WHIM. In the coming decade, our understanding of X-ray emissions from filaments and the WHIM within them will significantly improve from multiple aspects. The 4-metre Multi-Object Spectroscopic Telescope (4MOST) Cosmology Redshift Survey (Richard et al. 2019) will bring deep and wide field galaxy spectroscopic redshift data in the southern hemisphere, allowing us to take full use of the eRASS data of the German consortium. Improved filament finders such as the Monte Carlo Physarum Machine (Elek et al. 2022) and 1-DREAM (Canducci et al. 2022) will be applied to new survey data and provide a more accurately constructed filament catalog. Meanwhile, a better understanding of X-ray properties of galaxy groups, AGNs, and XRBs will help us to obtain a better estimate of the contamination fractions of unmasked sources. Moreover, high spectral resolution X-ray missions such as Hot Universe Baryon Surveyor (Cui et al. 2020) and Line Emission Mapper (Kraft et al. 2022) will be able to explore a wider parameter space of the WHIM properties.
@@ Acknowledgements.
We acknowledge Jelle Kaastra for the instruction on using the SPEX Pion model. We also acknowledge the anonymous referee for the insightful comments that improved the manuscript. X.Z., E.B., V.G., A.L., C.G., and S.Z. acknowledge financial support from the European Research Council (ERC) Consolidator Grant under the European Union’s Horizon 2020 research and innovation program (grant agreement CoG DarkQuest No 101002585). M.C.H.Y. and M.F. acknowledge support from the Deutsche Forschungsgemeinschaft through the grant FR 1691/2-1. Y.Z. and G.P. acknowledge funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 865637). G.P. acknowledges support from Bando per il Finanziamento della Ricerca Fondamentale 2022 dell’Istituto Nazionale di Astrofisica (INAF): GO Large program and from the Framework per l’Attrazione e il Rafforzamento delle Eccellenze (FARE) per la ricerca in Italia (R20L5S39T9). M.B. acknowledges funding by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy – EXC 2121 “Quantum Universe” – 390833306. A.V. acknowledges funding by the Deutsche Forschungsgemeinschaft – 450861021. N. M. acknowledges funding by the European Union through a Marie Skłodowska-Curie Action Postdoctoral Fellowship (Grant Agreement: 101061448, project: MEMORY). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or of the Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument were led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The eROSITA data shown here were processed using the eSASS/NRTA software system developed by the German eROSITA consortium.